OpenStudy (anonymous):

rectangular form 6(cos(150) + isin(150)) a.(3,√3,-3) B.(-3,√2,3) C.(3,√2,-3) D.(-3,√3,3)

OpenStudy (anonymous):

your job is to evaluate \(\cos(150)\) and \(\sin(150)\) then multiply the result by 6

OpenStudy (ryan123345):

Hey there! I see that you are new to openstudy. This is a great website on getting help with questions of any subject. It is a fun way of learning! I wish you good luck on openstudy and dont forget to read the http://openstudy.com/code-of-conduct so you know the rules and regulations :) Hope you have fun!

OpenStudy (anonymous):

none of the above are correct

OpenStudy (anonymous):

for a variety of reasons first, standard form is not ordered pairs, it is \(a+bi\) second, \(\cos(150)=-\frac{\sqrt{3}}{2}\) and \(\sin(150)=\frac{1}{2}\)

OpenStudy (anonymous):

the correct answer is \[-3\sqrt{3}+3i\]

OpenStudy (ryan123345):

i love your confidence @satellite73 lol

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

still none of the answers are right, because a complex number in rectangular form is \(a+bi\) not \((a,b)\) whoever wrote this is very confused

OpenStudy (anonymous):

i don't mean you, i mean the person who asked the question to begin with